# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v3

@@ 4 Numerical Experiments
In this section, we conduct numerical experiments to evaluate the performance and scalability of the proposed PG-DPO framework (both Baseline and Projected PG-DPO (P-PGDPO) variants). We focus on the multi-asset, multi-factor Ornstein-Uhlenbeck (OU) portfolio optimization problem introduced in Section 2.3.2. The objective is to maximize the expected utility of terminal wealth, $J=\mathbb{E}[U(X_{T})]$, using a CRRA utility function with risk aversion $\gamma=2.0$. For these experiments, we specialize the general model by assuming no intermediate consumption (i.e., $C_{t}\equiv 0$) and focus solely on the utility from terminal wealth (effectively setting the bequest weight $\kappa=1$ if $U(X_{T})$ is the direct objective, or adjusting $\kappa$ appropriately if a specific HARA utility form with $a\neq 0$ was being precisely mapped).
This specific affine model setting allows for comparison against a semi-analytical benchmark solution derived from the HJB equation (details in Appendix B.2). Furthermore, to provide a comprehensive assessment against other contemporary deep learning techniques for high-dimensional control problems, we also compare our PG-DPO framework with the Deep BSDE method (Han et al., , 2018; Weinan, , 2017). The Deep BSDE approach reformulates the stochastic control problem in terms of a system of backward stochastic differential equations, approximating the unknown components of the solution (typically related to the value function gradient or control variables) using neural networks. These networks are trained by minimizing the error incurred in satisfying the BSDE dynamics and terminal conditions through Monte Carlo simulation.
We compare the policies generated by Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and the Deep BSDE method against the HJB-based benchmark using the Root Mean Squared Error (RMSE) metric across various dimensions, scaling up to n=50 assets and k=10 state factors. For complete details on the specific parameter values, the implementations of PG-DPO and Deep BSDE (including network architectures and training setups), benchmark computation, and evaluation procedures, please refer to Appendix D. The subsequent subsections present and analyze the results of these comparative experiments.
@@ 4.1 Convergence Speed and Accuracy Analysis
Figure 1: Average policy RMSE comparison between Baseline PG-DPO (e.g., blue), Projected PG-DPO (P-PGDPO) (e.g., orange), and Deep BSDE (e.g., green) relative to the benchmark solution across different asset dimensions ($n$) with $k=10$ factors fixed. Values averaged over $k$ evaluation ‘slices.’ For each slice $j=1,\dots,k$, the evaluation (e.g., for RMSE calculation or visualization) focuses on state factor $Y_{j}$, while other factors $Y_{l},l\neq j,$ are held constant at their respective long-term means $\theta_{Y,l}$. Further details on the evaluation methodology are provided in Appendix D. Note the logarithmic scale on the $y$-axis.
This section quantitatively compares the convergence speed and final accuracy of the proposed PG-DPO framework (Baseline and Projected PG-DPO (P-PGDPO) variants) against the Deep BSDE method. We analyze their performance across various dimensional settings, focusing on how the number of assets ($n$) and the number of state factors ($k$) impact accuracy. Figure 1 (Note: Assuming you will update the figure label in your main document if it includes Deep BSDE results) displays the average Root Mean Squared Error (RMSE) between the learned/derived policies and the benchmark optimal policy. The RMSE is plotted against training epochs on a logarithmic scale for configurations with $k=10$ state variables and varying numbers of assets ($n=1,10,50$).
@@
The RMSE for P-PGDPO represents the error of the total portfolio policy $\bm{\pi}^{\mathrm{PMP}}$ (Eq. (11)), which combines myopic and hedging components. As detailed in Section 4.2, the myopic component’s error is typically negligible ($O(10^{-7})$). Thus, the total RMSE for P-PGDPO is almost entirely driven by the accuracy of estimating the hedging component using BPTT-derived costate derivatives, $\partial_{\mathbf{Y}}\lambda_{t}^{*}$.
Table 2 summarizes the minimum achieved policy RMSE and the epochs at which these minima were observed for the $k=10$ scenario with varying $n$. A consistent observation is the superior performance of the P-PGDPO method, which achieves significantly lower RMSE values compared to both Baseline PG-DPO and Deep BSDE across all tested asset dimensions. For instance, with $n=1,k=10$, P-PGDPO’s minimum RMSE is approximately an order of magnitude better than Deep BSDE and two orders of magnitude better than Baseline PG-DPO. Deep BSDE generally performs between the Baseline and P-PGDPO methods. While P-PGDPO maintains its lead as $n$ increases to 50, the performance gap between P-PGDPO and Deep BSDE tends to narrow, suggesting that the increasing dimensionality of $n$ poses challenges even for the PMP-based analytical solution, possibly due to the ill-conditioning of the covariance matrix $\Sigma$. All methods show an increase in RMSE as $n$ grows, indicating the inherent difficulty of high-dimensional asset allocation.
@@
Table 2: Summary of Minimum Policy RMSE and Convergence Epochs
Dimension Method Min. Policy RMSE Epochs at Min.
($n$ assets, $k=10$ factors)
n=1, k=10 Baseline PG-DPO $1.75\times 10^{-2}$ 2800
P-PGDPO $4.72\times 10^{-4}$ 1200
Deep BSDE $4.78\times 10^{-3}$ 600
n=10, k=10 Baseline PG-DPO $5.11\times 10^{-2}$ 9600
P-PGDPO $5.26\times 10^{-3}$ 1000
Deep BSDE $3.46\times 10^{-2}$ 1400
n=50, k=10 Baseline PG-DPO $1.05\times 10^{-1}$ 8800
P-PGDPO $1.58\times 10^{-2}$ 5400
Deep BSDE $4.90\times 10^{-2}$ 9200
Note: Minimum Policy RMSE values are averaged across $k$ evaluation ’slices.’ For each slice $j=1,\dots,k$, the evaluation (e.g., for RMSE calculation) focuses on state factor $Y_{j}$, while other factors $Y_{l},l\neq j,$ are held constant at their respective long-term means $\theta_{Y,l}$. Further details on the evaluation methodology are provided in Appendix D. Epochs indicate the recorded iteration number when the minimum average RMSE was observed; evaluations and logging were performed every 200 iterations.
First, Table 3 presents results where the number of assets is fixed at $n=1$, and the number of state factors $k$ is varied ($k=1,5,10$). In this scenario, the performance of all three methodologies remains relatively stable across different values of $k$. For P-PGDPO, the RMSE even shows a slight improvement as $k$ increases, possibly due to more information being available for costate estimation. Deep BSDE and Baseline PG-DPO also exhibit no significant degradation in performance. This suggests that, for a small number of assets, the ”curse of dimensionality” associated with an increasing number of state factors ($k$) is largely managed by these neural network-based approaches. The underlying PMP structure or BSDE formulation appears robust to the increase in $k$ when $n$ is small and fixed, with Monte Carlo sampling and function approximation being the primary challenges rather than ill-conditioned linear algebra.
@@
Table 3: Summary of Minimum Policy RMSE and Convergence Epochs for $n=1$ (Varying $k$)
$k$ factors Method Min. Policy RMSE Epochs at Min.
@@
1 Baseline PG-DPO $3.26\times 10^{-2}$ 1000
P-PGDPO $1.20\times 10^{-3}$ 800
Deep BSDE $7.39\times 10^{-3}$ 800
5 Baseline PG-DPO $3.71\times 10^{-2}$ 1400
P-PGDPO $2.64\times 10^{-4}$ 800
Deep BSDE $1.02\times 10^{-2}$ 1400
10 Baseline PG-DPO $1.75\times 10^{-2}$ 2800
P-PGDPO $4.72\times 10^{-4}$ 1200
Deep BSDE $4.78\times 10^{-3}$ 600
Table 4: Summary of Minimum Policy RMSE and Convergence Epochs for $k=2$ (Varying $n$)
$n$ assets Method Min. Policy RMSE Epochs at Min.
@@
1 Baseline PG-DPO $4.77\times 10^{-2}$ 1200
P-PGDPO $1.98\times 10^{-4}$ 800
Deep BSDE $4.63\times 10^{-3}$ 3400
10 Baseline PG-DPO $1.12\times 10^{-1}$ 4000
P-PGDPO $1.22\times 10^{-2}$ 1400
Deep BSDE $2.32\times 10^{-2}$ 2600
50 Baseline PG-DPO $2.24\times 10^{-1}$ 7200
P-PGDPO $5.55\times 10^{-2}$ 4400
Deep BSDE $1.88\times 10^{-1}$ 4400
Second, Table 4 shows results where the number of state factors is fixed at $k=2$, and the number of assets $n$ is varied ($n=1,10,50$). In stark contrast to the previous case, increasing $n$ leads to a significant deterioration in performance for all three methods. Baseline PG-DPO shows the most substantial increase in RMSE. While P-PGDPO still yields the lowest RMSE, its error also increases considerably with $n$. Deep BSDE’s performance also degrades, approaching that of Baseline PG-DPO for $n=50$. This pronounced degradation with increasing $n$ (even for a small $k$) points towards a different kind of challenge not solely related to the dimensionality of the state space for function approximation. A primary suspect is the numerical instability arising from the asset covariance matrix $\Sigma$. As $n$ increases, $\Sigma$ ($n\times n$) is more likely to become ill-conditioned (i.e., have a high condition number). The optimal policy calculation, particularly in the PMP framework (Eq. (11)), involves the inverse $\Sigma^{-1}$. If $\Sigma$ is ill-conditioned, any small errors in the estimated costates or market parameters can be greatly amplified by $\Sigma^{-1}$, leading to large errors in the computed policy. This suggests that the “curse” associated with large $n$ is not just about learning a higher-dimensional policy function but also about the increasing fragility of the underlying mathematical problem structure due to the properties of $\Sigma$. The problem itself becomes “damaged” or numerically challenging, impacting any method that relies on or implicitly reflects this structure.
These findings underscore that while neural network approaches have made significant strides in handling high-dimensional state spaces (related to $k$), the challenges posed by a large number of decision variables (assets $n$), especially when intertwined with ill-conditioned covariance matrices, require further attention. Strategies such as covariance matrix shrinkage or other regularization techniques applied prior to or within the optimization might be necessary to effectively address this “asset-dimension curse”.
@@ 4.2 High-Dimensional Performance Analysis ($n=50,k=10$)
We now visually examine the performance of the three methodologies in the most challenging setting considered: $n=50$ assets and $k=10$ state factors. This high-dimensional scenario, where traditional DP methods are intractable, serves to highlight the distinct capabilities and limitations of each approach. As a reference, Figure 2 displays the benchmark analytical solution for the first asset ($\pi_{1}^{\text{sol}}$), decomposed into its myopic ($\pi_{1,0}^{\text{sol}}$) and intertemporal hedging ($\pi_{1,1}^{\text{sol}}$) components. The hedging component is clearly non-trivial and exhibits complex dependencies on both the state variable $Y_{1}$ and time-to-maturity, underscoring its importance for optimal allocation.
@@
Figure 3 provides a side-by-side comparison of the policies learned by Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and Deep BSDE, alongside their respective errors, for the first asset. These plots correspond to iterations where each method achieved its (near) optimal performance in the $n=50,k=10$ setting, as indicated in Table 2 (e.g., iteration 5400 for P-PGDPO, which is the nearest logged checkpoint to the observed minimum RMSE at iteration 5399, as evaluations were performed every 200 iterations).
@@
Figure 3: Comparison of total policies (left column) and policy errors (right column) for the first asset ($i=1$) across Baseline PG-DPO (top row), Projected PG-DPO (P-PGDPO) (middle row), and Deep BSDE (bottom row) in the $n=50,k=10$ case. Results shown at (near) optimal iterations for each method (e.g., iteration 5400 for P-PGDPO), plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (consistent with Figure 2). Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$. Errors use a symmetric logarithmic (symlog) scale. Different color bar ranges reflect varying policy/error magnitudes.
The Baseline PG-DPO method (Figure 3a,b), despite utilizing a standard neural network architecture (detailed in Appendix D.2), struggles significantly. While its learned policy exhibits some non-linear structure, it bears little resemblance to the benchmark, and the corresponding error is extremely large. This performance visually confirms the poor quantitative results in Table 2 and highlights the difficulty of directly approximating the high-dimensional optimal policy function using a neural network without leveraging more structural information from the problem.
@@
Figure 4: Decomposition of the Projected PG-DPO (P-PGDPO) policy for the first asset ($i=1$) in the $n=50,k=10$ case (results from iteration 5400), plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (consistent with Figure 2). Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$. Top row: myopic and hedging components. Bottom row: their respective errors (symmetric logarithmic (symlog) scale). Note different color bar ranges.
The Deep BSDE method (Figure 3e,f) shows a marked improvement over the Baseline. Its learned policy qualitatively resembles the myopic component of the benchmark (cf. Figure 2b), primarily capturing dependencies on the state variable $Y_{1}$. However, the policy largely fails to capture the complex intertemporal hedging demands. The error plot for Deep BSDE (Figure 3f) shows significant discrepancies in regions where hedging is important, suggesting that the hedging component is largely absent from the Deep BSDE solution. This difficulty in capturing hedging demands can be attributed to several factors. The Deep BSDE method primarily focuses on matching the terminal condition and ensuring consistency of the value process $Y_{t}$ and the gradient-related process $Z_{t}$ (approximating first-order derivatives of the value function) through its loss function. Intertemporal hedging, however, critically depends on the mixed second-order derivative $V_{x\mathbf{y}}$, which is not explicitly represented or directly targeted by the standard two-network Deep BSDE architecture. Furthermore, the loss function penalizes inaccuracies in $Z_{t}$ primarily through a quadratic martingale term. Consequently, small errors in estimating the gradient components relevant to $V_{x\mathbf{y}}$ might have a limited impact on the overall training objective but can be significantly magnified when constructing the policy, especially after multiplication by the potentially ill-conditioned inverse covariance matrix $\Sigma^{-1}$ in high-dimensional settings. As a result, while the Deep BSDE approach can effectively learn the dominant myopic component of the policy, it may struggle to precisely fit the more nuanced intertemporal hedging patterns.
@@
In stark contrast, the Projected PG-DPO (P-PGDPO) method (Figure 3c,d) delivers a policy that is visually almost identical to the benchmark total policy (Figure 2a). The corresponding error is substantially smaller than that of the other two methods. This superior performance underscores the efficacy of its decoupled learning structure: first estimating costates via BPTT and stabilizing them through Monte Carlo averaging, then analytically constructing the policy using the PMP FOCs.
@@
The remarkable accuracy of P-PGDPO is further elucidated by decomposing its policy into myopic and hedging components, as shown in Figure 4 (results from iteration 5400). The myopic component (Figure 4(a)) is virtually identical to the benchmark myopic component, with its error (Figure 4(c)) being negligible ($O(10^{-7})$). The hedging component derived by P-PGDPO (Figure 4(b)) also successfully captures the complex shape of the benchmark hedging demand (Figure 2c). The remaining error in the hedging component (Figure 4(d)) dictates the overall error of the P-PGDPO policy. This residual hedging error is likely not due to a fundamental flaw in the P-PGDPO algorithm itself but rather a consequence of the numerical challenges inherent in high-dimensional asset allocation. Specifically, the hedging term in the PMP FOC involves $\Sigma^{-1}$:
@@
$\pi^{\text{hedge}}=-\frac{1}{X_{t}V_{xx}}\,\underbrace{\Sigma^{-1}}_{\text{cond num can be large for }n\uparrow}(\bm{\sigma}\,\bm{\rho}\,\bm{\sigma}_{Y})\,V_{x\mathbf{Y}}.$

### v4

@@ 4 Numerical Experiments
In this section, we conduct numerical experiments to evaluate the performance and scalability of the proposed PG-DPO framework, including both the baseline and Projected PG-DPO (P-PGDPO) variants. The experiments are based on the multi-asset, multi-factor Ornstein–Uhlenbeck (OU) portfolio optimization problem introduced in Section 2.3.2. For this setting, we specialize the general model by assuming no intermediate consumption (i.e., $C_{t}=0$) and focus exclusively on utility from terminal wealth, with $\kappa=1$, so that the objective becomes $J=\mathbb{E}[U(X_{T})]$.
This affine setting admits a semi-analytical benchmark solution, derived from the associated HJB equation, which we use as a reference for evaluation (see Appendix B.2). To further assess the effectiveness of our approach, we compare PG-DPO with a state-of-the-art deep learning method for high-dimensional control problems—the Deep BSDE framework (Han et al., , 2018; Weinan, , 2017).
The Deep BSDE method reformulates the stochastic control problem as a system of backward stochastic differential equations (BSDEs). Neural networks are trained to approximate key solution components, typically related to the value function and its gradient. These approximations are then used to construct the control policy, often by enforcing or approximating first-order optimality conditions derived from the Hamilton–Jacobi–Bellman equation. The networks are trained by minimizing residuals in the BSDE dynamics and terminal condition using Monte Carlo simulation.
We evaluate the control policies produced by Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and the Deep BSDE method (when used for control) against the semi-analytical benchmark. The evaluation metric is Root Mean Squared Error (RMSE) in portfolio weights, assessed across a range of problem sizes, including up to $n=50$ assets and $k=10$ state factors.
Complete details on model parameters, PG-DPO and Deep BSDE implementations (including neural network architectures and training protocols), benchmark computation, and evaluation methodology are provided in Appendix D. The following subsections present and analyze the results of these comparative experiments.
@@ 4.1 Convergence Speed and Accuracy Analysis
Figure 1: Average policy RMSE comparison between Baseline PG-DPO (e.g., blue), Projected PG-DPO (P-PGDPO) (e.g., orange), and Deep BSDE (e.g., green) relative to the benchmark solution across different asset dimensions ($n$) with $k=10$ factors. Values averaged over $k=10$ evaluation ‘slices.’ For each slice $j=1,\dots,10$, the evaluation (e.g., for RMSE calculation or visualization) focuses on state factor $Y_{j}$, while other factors $Y_{l},l\neq j,$ are held constant at their respective long-term means $\theta_{Y,l}$. Further details on the evaluation methodology are provided in Appendix D. Note the logarithmic scale on the $y$-axis.
This section quantitatively compares the convergence speed and final accuracy of the proposed PG-DPO framework—including both the Baseline and Projected PG-DPO (P-PGDPO) variants—against the Deep BSDE method, when applied to derive control policies. We evaluate performance across various problem dimensions, focusing on how the number of assets ($n$) and the number of state factors ($k$) affect policy accuracy.
Figure 1 shows the average Root Mean Squared Error (RMSE) between the learned or derived policies and the semi-analytical benchmark policy. The RMSE is plotted against the number of training iterations on a logarithmic scale for experiments with $k=10$ state variables and varying asset counts ($n=1,10,50$).
@@
For P-PGDPO, the reported RMSE corresponds to the total portfolio policy $\bm{\pi}^{\mathrm{PMP}}$ (Eq. (11)), which combines the myopic and intertemporal hedging components. As discussed in Section 4.2, the myopic error is typically negligible—on the order of $O(10^{-7})$. Thus, the total RMSE for P-PGDPO is effectively governed by the accuracy of the hedging component, which depends on estimating the costate gradient $\partial_{\mathbf{Y}}\lambda_{t}^{*}$ via BPTT.
Table 2: Summary of Minimum Policy RMSE and Convergence Iterations
Dimension Method Min. Policy RMSE Iterations at Min.
($n$ assets, $k=10$ factors)
n=1, k=10 Baseline PG-DPO $1.751\times 10^{-2}$ 2800
P-PGDPO $4.710\times 10^{-4}$ 1200
Deep BSDE $8.146\times 10^{-2}$ 1
n=10, k=10 Baseline PG-DPO $4.996\times 10^{-2}$ 9800
P-PGDPO $2.799\times 10^{-3}$ 4800
Deep BSDE $7.623\times 10^{-2}$ 200
n=50, k=10 Baseline PG-DPO $1.070\times 10^{-1}$ 8800
P-PGDPO $1.522\times 10^{-2}$ 10000
Deep BSDE $8.199\times 10^{-2}$ 1200
Note: Minimum policy RMSE values are averaged across $k$ evaluation slices. For each slice $j=1,\dots,k$, the evaluation (e.g., for RMSE calculation) focuses on state factor $Y_{j}$, while all other factors $Y_{\ell}$, for $\ell\neq j$, are fixed at their respective long-run means $\theta_{Y,\ell}$. Further details on the evaluation procedure are provided in Appendix D. The iteration numbers indicate when the minimum average RMSE was observed; evaluations and logging were performed every 200 iterations.
Table 2 summarizes the minimum achieved policy RMSE values and the corresponding iteration counts at which these minima occur, across $k=10$ evaluation ‘slcies’ with varying numbers of assets $n$. In each evaluation slice $j=1,\dots,10$, the RMSE and visualization are computed with respect to state factor $Y_{j}$, while all other factors $Y_{\ell}$, for $\ell\neq j$, are fixed at their respective long-term means $\theta_{Y,\ell}$. Further details on the evaluation methodology are provided in Appendix D.
A consistent and notable finding is the superior performance of the P-PGDPO method. It achieves substantially lower RMSEs than both Baseline PG-DPO and Deep BSDE across all tested asset dimensions. For example, in the $n=1$ case, P-PGDPO attains a minimum RMSE of $4.710\times 10^{-4}$, significantly outperforming Baseline PG-DPO ($1.751\times 10^{-2}$) and Deep BSDE ($8.146\times 10^{-2}$). This result highlights the effectiveness of P-PGDPO in delivering high-accuracy control policies.
The comparison between Baseline PG-DPO and Deep BSDE is more nuanced. Deep BSDE often reaches its minimum RMSE in fewer iterations—e.g., 1 iteration for $n=1$, and 200 iterations for $n=10$—compared to 2800 and 9800 iterations, respectively, for Baseline PG-DPO (see Table 2). This suggests faster initial convergence for Deep BSDE. However, its final RMSE is not consistently lower than that of Baseline PG-DPO. For both $n=1$ and $n=10$, Baseline PG-DPO ultimately achieves better accuracy: $1.751\times 10^{-2}$ vs. $8.146\times 10^{-2}$, and $4.996\times 10^{-2}$ vs. $7.623\times 10^{-2}$, respectively.
In the high-dimensional case of $n=50$, however, Deep BSDE achieves a lower final RMSE ($8.199\times 10^{-2}$) than Baseline PG-DPO ($1.070\times 10^{-1}$). As illustrated in Figure 1, after early convergence, Deep BSDE’s RMSE may plateau or exhibit greater variance, while Baseline PG-DPO often demonstrates more gradual but continued improvement in certain configurations.
P-PGDPO maintains a significant accuracy advantage even as $n$ increases to 50. However, all methods exhibit rising RMSE with increasing $n$ (see Table 2), reflecting the inherent challenges of high-dimensional asset allocation—potentially exacerbated by issues such as the ill-conditioning of the covariance matrix $\Sigma$, as discussed later.
@@
First, Table 3 reports results for the case where the number of assets is fixed at $n=1$, and the number of state factors is varied ($k=1,5,10$). In this setting, P-PGDPO consistently achieves the lowest RMSE across all configurations. Specifically, its RMSE is $1.120\times 10^{-3}$ for $k=1$, improves to $2.500\times 10^{-4}$ for $k=5$, and is $4.710\times 10^{-4}$ for $k=10$. While not strictly monotonic, P-PGDPO demonstrates strong performance across all $k$, with the best result observed at $k=5$.
The performances of Baseline PG-DPO and Deep BSDE are more varied. For $n=1$:
  • At $k=1$: Deep BSDE ($9.350\times 10^{-3}$) outperforms Baseline PG-DPO ($3.260\times 10^{-2}$).
  • At $k=5$: Deep BSDE ($3.094\times 10^{-2}$) outperforms Baseline PG-DPO ($3.714\times 10^{-2}$).
  • At $k=10$: Baseline PG-DPO ($1.751\times 10^{-2}$) outperforms Deep BSDE ($8.146\times 10^{-2}$); however, Deep BSDE converges to its minimum in just one iteration.
These results suggest that, when the number of assets is small, the “curse of dimensionality” associated with increasing $k$ is largely mitigated by neural network–based approaches. Nevertheless, their relative performance can shift. Both the PMP and BSDE formulations appear robust to increases in factor dimensionality when $n$ is fixed and small. In this case, Monte Carlo sampling error and function approximation quality become the primary computational bottlenecks, rather than issues related to ill-conditioned linear algebra.
Second, Table 4 reports results where the number of state factors is fixed at $k=2$, and the number of assets $n$ is varied ($n=1,10,50$). In stark contrast to the previous case, increasing $n$ leads to a substantial deterioration in performance across all three methods.
Baseline PG-DPO shows the most pronounced increase in RMSE, rising from $4.770\times 10^{-2}$ for $n=1$ to $3.875\times 10^{-1}$ for $n=50$. While P-PGDPO still achieves the lowest RMSE in all cases, its error also increases with $n$—from $1.990\times 10^{-4}$ to $7.330\times 10^{-2}$. Deep BSDE’s performance degrades most dramatically, with RMSE rising from $2.675\times 10^{-3}$ at $n=1$ to $6.340\times 10^{-1}$ at $n=50$. Notably, in the $n=50,k=2$ case, Baseline PG-DPO ($3.875\times 10^{-1}$) outperforms Deep BSDE ($6.340\times 10^{-1}$).
This pronounced degradation with increasing $n$, even for small $k$, points to a distinct source of difficulty—one not solely attributable to function approximation in high-dimensional state spaces. A likely culprit is the numerical instability introduced by the asset covariance matrix $\Sigma$. As $n$ increases, the $n\times n$ matrix $\Sigma$ becomes more prone to ill-conditioning (i.e., a high condition number). Since optimal portfolio policies in the PMP framework (Eq. (11)) involve computing $\Sigma^{-1}$, small estimation errors in the costates or model parameters can be significantly amplified, resulting in large policy errors.
This observation suggests that the “curse” associated with large $n$ is not merely about learning a higher-dimensional function, but also reflects the increasing fragility of the optimization problem itself. The structure of the problem becomes numerically unstable, which adversely affects any method that explicitly or implicitly depends on $\Sigma$ or its inverse.
These findings highlight that while neural network–based methods have made substantial progress in managing high-dimensional state spaces (associated with large $k$), challenges related to a large number of decision variables (i.e., assets $n$)—particularly in the presence of ill-conditioned covariance matrices—remain a significant concern. Addressing this “asset-dimension curse” may require incorporating techniques such as covariance matrix shrinkage, spectral regularization, or other forms of preconditioning within the optimization problem.
@@
Table 3: Summary of Minimum Policy RMSE and Convergence Iterations for $n=1$ (Varying $k$)
$k$ factors Method Min. Policy RMSE Iterations at Min.
@@
1 Baseline PG-DPO $3.260\times 10^{-2}$ 1000
P-PGDPO $1.120\times 10^{-3}$ 800
Deep BSDE $9.350\times 10^{-3}$ 3600
5 Baseline PG-DPO $3.714\times 10^{-2}$ 1400
P-PGDPO $2.500\times 10^{-4}$ 800
Deep BSDE $3.094\times 10^{-2}$ 9400
10 Baseline PG-DPO $1.751\times 10^{-2}$ 2800
P-PGDPO $4.710\times 10^{-4}$ 1200
Deep BSDE $8.146\times 10^{-2}$ 1
Table 4: Summary of Minimum Policy RMSE and Convergence Iterations for $k=2$ (Varying $n$)
$n$ assets Method Min. Policy RMSE Iterations at Min.
@@
1 Baseline PG-DPO $4.770\times 10^{-2}$ 1200
P-PGDPO $1.990\times 10^{-4}$ 800
Deep BSDE $2.675\times 10^{-3}$ 3600
10 Baseline PG-DPO $1.145\times 10^{-1}$ 4000
P-PGDPO $3.645\times 10^{-3}$ 4600
Deep BSDE $3.835\times 10^{-2}$ 1
50 Baseline PG-DPO $3.875\times 10^{-1}$ 5600
P-PGDPO $7.330\times 10^{-2}$ 9800
Deep BSDE $6.340\times 10^{-1}$ 3800
@@ 4.2 High-Dimensional Performance Analysis ($n=50,k=10$)
We now visually examine the performance of the three methodologies in the most challenging setting considered: $n=50$ assets and $k=10$ state factors. This high-dimensional configuration, where traditional dynamic programming methods are computationally intractable, highlights the distinct capabilities and limitations of each approach.
As a reference, Figure 2 displays the benchmark analytical solution for the first asset ($\pi_{1}^{\text{sol}}$), decomposed into its myopic ($\pi_{1,0}^{\text{sol}}$) and intertemporal hedging ($\pi_{1,1}^{\text{sol}}$) components. The hedging component is clearly non-trivial, exhibiting complex dependencies on both the state variable $Y_{1}$ and the remaining time to maturity. This underscores its importance for achieving optimal portfolio allocation in dynamic settings.
@@
Figure 3 presents a side-by-side comparison of the policies learned by Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and Deep BSDE, along with their corresponding errors, for the first asset. These plots correspond to the iterations at which each method achieved its (near) optimal performance in the $n=50,k=10$ setting, as reported in Table 2. For instance, the policy shown for P-PGDPO corresponds to iteration 6400, which is the nearest logged evaluation point to the observed minimum RMSE at iteration 5399 (evaluation was performed every 200 iterations, and the global minimum RMSE for P-PGDPO in this setting was ultimately achieved at iteration 10000, as also noted in Table 2).
@@
Figure 3: Comparison of total policies (left column) and policy errors (right column) for the first asset ($i=1$) across Baseline PG-DPO (top row), Projected PG-DPO (P-PGDPO) (middle row), and Deep BSDE (bottom row) in the $n=50,k=10$ case. Results shown at (near) optimal iterations for each method (e.g., iteration 6400 for P-PGDPO as plotted), plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (consistent with Figure 2). Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$. Errors use a symmetric logarithmic (symlog) scale. Different color bar ranges reflect varying policy/error magnitudes.
The Baseline PG-DPO method (Figure 3a,b), employing a standard neural network architecture (see Appendix D.2), performs poorly in this high-dimensional setting. Although the learned policy exhibits some non-linear structure, it bears little resemblance to the benchmark solution, and the corresponding error is substantial. This outcome visually corroborates the relatively high RMSE of $1.070\times 10^{-1}$ reported for Baseline PG-DPO in the $n=50,k=10$ case (Table 2), and underscores the difficulty of directly approximating a high-dimensional optimal policy function using generic neural architectures without incorporating structural insights from the problem’s analytic formulation.
@@
Figure 4: Decomposition of the Projected PG-DPO (P-PGDPO) policy for the first asset ($i=1$) in the $n=50,k=10$ case (results from iteration 6400 as plotted), plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (consistent with Figure 2). Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$. Top row: myopic and hedging components. Bottom row: their respective errors (symmetric logarithmic (symlog) scale). Note different color bar ranges.
The Deep BSDE method (Figure 3e,f) shows a marked improvement over Baseline PG-DPO. Its learned policy qualitatively resembles the myopic component of the benchmark (cf. Figure 2b), primarily capturing the dependency on the state variable $Y_{1}$. However, the policy fails to capture the more complex intertemporal hedging demands.
The corresponding error plot (Figure 3f) reveals significant discrepancies in regions where hedging plays a critical role, suggesting that the hedging component is largely absent from the Deep BSDE solution. This shortcoming can be attributed to both architectural and training limitations.
The Deep BSDE framework is primarily designed to match the terminal condition and ensure consistency of the forward value process $Y_{t}$ (approximating the value function) and the backward process $Z_{t}$ (approximating its gradient). However, intertemporal hedging depends crucially on the mixed second-order derivative $V_{x\mathbf{y}}$, which is neither explicitly represented nor directly supervised in the standard two-network Deep BSDE setup.
Moreover, the loss function penalizes deviations in $Z_{t}$ through a quadratic martingale residual, which may assign relatively low weight to errors in specific gradient directions, particularly those influencing $V_{x\mathbf{y}}$. These small gradient estimation errors can nonetheless be magnified when constructing the control policy, especially through multiplication by the potentially ill-conditioned inverse covariance matrix $\Sigma^{-1}$ in high-dimensional settings.
As a result, while the Deep BSDE method effectively learns the dominant myopic component of the policy, it tends to underperform in capturing the subtler hedging demands essential for dynamic portfolio optimization.
@@
In stark contrast, the Projected PG-DPO (P-PGDPO) method (Figure 3c,d) produces a policy that is visually nearly indistinguishable from the benchmark total policy (Figure 2a). The corresponding error is substantially smaller than that of the other two methods, consistent with its superior RMSE of $1.522\times 10^{-2}$ for the $n=50,k=10$ case (Table 2). This performance underscores the effectiveness of P-PGDPO’s decoupled learning structure: first estimating costates via BPTT and stabilizing them through Monte Carlo averaging, then analytically constructing the policy using the PMP first-order conditions.
The exceptional accuracy of P-PGDPO is further illuminated by decomposing its policy into myopic and hedging components, as shown in Figure 4 (from iteration 6400). The myopic component (Figure 4(a)) closely matches the benchmark, with its error (Figure 4(c)) being negligible—on the order of $O(10^{-7})$.
@@
The hedging component (Figure 4(b)) also successfully captures the complex shape of the benchmark hedging demand (Figure 2c). The residual error (Figure 4(d)) in this component dominates the total policy error. Notably, this hedging error likely does not stem from a flaw in the P-PGDPO algorithm itself, but rather reflects the numerical challenges inherent to high-dimensional asset allocation. Specifically, the hedging term in the PMP solution involves the inverse of the asset return covariance matrix $\Sigma$:
@@
$\pi^{\text{hedge}}=-\frac{1}{X_{t}V_{xx}}\,\underbrace{\Sigma^{-1}}_{\text{canbe ill-conditioned for large }n}(\bm{\sigma}\,\bm{\rho}\,\bm{\sigma}_{Y})\,V_{x\mathbf{Y}}.$